The elastic limit of sgeel is 5.0Pa , find the maximum diameter of steel wire that can support a 700N weight without exceeding its elasgic limit.


1
Expert's answer
2021-12-06T09:41:40-0500
"\frac{700}{5}=0.25\pi d^2\\d=13.4\ m"